find the vertex of parabola x²=2(2x+y).just give me final answer

Grade:11

1 Answers

Pradeep Tangirala
6 years ago
Vertex is (2,-2)...
x*2=4x+2y
x*2 - 4x = 2y
x*2 - 4x + 4 - 4 = 2y
(x-2)*2 = 2y + 4 
So....(x-2)*2 = 2(y+2)
So....vertex is....(2,-2)